# Dual Caption Preference Optimization for Diffusion Models

### v1

@@ Acknowledgments
Appendix A Background
A.1 Diffusion Models
Based on samples from a data distribution $q(x_{0})$, a noise scheduling function $\alpha_{t}$ and $\sigma_{t}$ (Rombach et al., 2022) denoising diffusion models (Song et al., 2020b) are generative models $p_{\theta}(x_{0})$ that operate through a discrete-time reverse process structured as a Markov Decision Proces where
$p_{\theta}(x_{t-1}|x_{t})=\mathcal{N}(x_{t-1};\mu_{\theta}(x_{t}),\sigma^{2}_{t|t-1}\frac{\sigma^{2}_{t-1}}{\sigma^{2}_{t}}I).$ (3)
The training process involves minimizing the evidence lower bound (ELBO) associated with this model (Song et al., 2021):
$L_{DM}=\mathit{\mathbb{E}}_{x_{0},\epsilon,t,x_{t}}[\omega(\lambda_{t})||\epsilon-\epsilon_{\theta}(x_{t},t)||^{2}_{2}]$ (4)
where $\epsilon\sim\mathcal{N}(0,I)$, $t\sim\mathcal{U}(o,T)$, $x_{t}~{}q(x_{t}|x_{0})=\mathcal{N}(x_{t};\alpha_{t}x_{0},\sigma^{2}_{t}I).\lambda_{t}=\alpha^{2}_{t}/\sigma^{2}_{t}$ is a signal-to-noise ratio (Kingma et al., 2021), $\omega(\lambda_{t})$ is a predefined weighting function (Song & Ermon, 2019).
A.2 Preference Optimization
$p_{BT}(x^{w}_{0}\succ x^{l}_{0}|c)=\sigma(r(c,x^{w}_{0})-r(c,x^{l}_{0}))$ (5)
where $\sigma$ denotes the sigmoid function, and $r(x_{0},c)$ is derived from a neural network parameterized by $\phi$, which is estimated through maximum likelihood training for binary classification as follows:
where the prompt $c$ and data pairs $x_{0}^{w}$, $x_{0}^{l}$ are sourced from a dataset that humans have annotated.
where $\beta$ controls how far the policy model $p_{\theta}$ can deviate from the reference model $p_{ref}$. It can be demonstrated that the objective in Equation 7 converges to the following policy model:
where $Z$ is the partition function, the training objective for $p_{\theta}$, inspired by DPO, has been derived to be equivalent to Equation 8 without the need for an explicit reward model $r(x,c)$. Instead, it learns directly from the preference data $(c,x^{w}_{0},x^{l}_{0})\sim D$.
$L_{\text{DPO}}(\theta)=-\mathit{\mathbb{E}}_{c,x^{w}_{0},x^{l}_{0}}[\log\sigma(\beta\log\frac{p_{\theta}(x^{w}_{0}|c)}{p_{\text{ref}}(x^{w}_{0}|c)}-\beta\log\frac{p_{\theta}(x^{l}_{0}|c)}{p_{\text{ref}}(x^{l}_{0}|c)})]$ (9)
where $\sigma$ represents the sigmoid function.
Through this reparameterization, instead of first optimizing the reward function $r$ and then applying reinforcement learning, the method directly optimizes the conditional distribution $p_{\theta}(x_{0}|c)$.
Appendix B Formal Proofs Regarding DCPO
B.1 Optimizing the DCPO Loss is Optimizing the DPO Loss
@@
$\begin{split}\mathcal{L}_{\text{DCPO}}(\theta;\mathcal{D^{\prime}},\beta,p_{\text{ref}})&=\mathit{\mathbb{E}}_{(x^{w}_{0},x^{l}_{0},z^{w},z^{l})\sim\mathcal{D^{\prime}}}\\ &\left[\log\left(\sigma\left(\beta\log\frac{p_{\theta}(x^{w}_{0},z^{w})}{p_{\text{ref}}(x^{w}_{0},z^{w})}-\beta\log\frac{p_{\theta}(x^{l}_{0},z^{l})}{p_{\text{ref}}(x^{l}_{0},z^{l})}\right)\right)\right]\\ &=\mathit{\mathbb{E}}_{(x^{w}_{0},x^{l}_{0},z^{w},z^{l})\sim\mathcal{D^{\prime}}}\\ &\left[\log\left(\sigma\left(\beta\log\frac{p_{\theta}(x^{w}_{0}|z^{w})p_{\theta}(z^{w})}{p_{\text{ref}}(x^{w}_{0}|z^{w})p_{\text{ref}}(z^{w})}\right.\right.\right.\left.\left.\left.-\beta\log\frac{p_{\theta}(x^{l}_{0}|z^{l})p_{\theta}(z^{l})}{p_{\text{ref}}(x^{l}_{0}|z^{l})p_{\text{ref}}(z^{l})}\right)\right)\right]\\ \end{split}$ (11)
@@
B.2 Analyses of DCPO’s Effectiveness
@@
$\Delta R_{\theta}\gg\Delta R_{\text{ref}}\implies\Delta_{\text{preferred}}\text{ increases.}$.
@@
Relative $\text{Relative Rate}=\frac{\Delta R_{\theta}}{\Delta R_{\text{ref}}}\gg 1,$
@@
Appendix C More Insights on DCPO
@@
Appendix D Pick-Double Caption Dataset
@@
Figure 9: Token distribution of original prompt.
We observed that some prompts contained only one or two words, while others were excessively long. To ensure a fair comparison, we removed prompts that were too short or too long, leaving us with approximately 17,000 instances. We then generated captions using two state-of-the-art models, LLaVA-1.6-34B, and Emu2-32B. Figure 10 provides examples from the dataset.
@@
Table 6 presents a statistical analysis of the Pick-Double Caption dataset. With the non-conditional prompt method, we found that the average token length of captions generated by LLaVA is similar to that of the original prompts. However, captions generated by LLaVA using conditional prompts are twice as long as the original prompts. Additionally, Emu2 generated captions that, on average, are half the length of the original prompts for both methods.
Table 6: Statistical information on the Pick-Double Caption dataset, including the CLIPscore of in-distribution data and average token count of captions generated by LLaVA and Emu2 for both in-distribution and out-of-distribution data.
[…]
prompt $c$ 15.95 15.95 (26.74, 25.41) (26.74, 25.41)
caption $z^{w}$ (LLaVA) 32.32 17.69 30.85 29.04
caption $z^{l}$ (LLaVA) 32.83 17.91 26.48 28.29
caption $z^{w}$ (Emu2) 7.75 8.40 25.44 25.18
caption $z^{l}$ (Emu2) 7.84 8.44 22.64 24.88
@@
Appendix E More Details on Perturbation Method
Table 7: Examples of perturbed prompts and captions after applying different levels of perturbation.
Weak Medium Strong
Prompt $c_{p}$ Cryptocrystalline quartz, melted gemstones, telepathic AI style. Painting of cryptocrystalline quartz. Melted gems. Sacred geometry. Cryptocrystalline quartz with melted stones, in telepathic AI style.
[…]
Original Prompt $c$: Painting of cryptocrystalline quartz melted gemstones sacred geometry pattern telepathic AI style
We provide the setups for the LLM-based perturbation process involved in the DCPO-p and DCPO-h pipelines. Similarly to the method of constructing paraphrasing adversarial attacks as synonym-swapping perturbation by Krishna et al. (2024), we use DIPPER (Krishna et al., 2024), a text generation model built by fine-tuning T5-XXL (Chung et al., 2022), to create semantically perturbed captions or prompts, as shown in Table 7. Our three levels of perturbation are achieved by only altering the setting of lexicon diversity (0 to 100) in DIPPER - we use 40 for Weak, 60 for Medium, and 80 for Strong. We also use ”Text perturbation for variable text-to-image prompt.” to prompt the perturbation. We hereby provide a code snippet to showcase the whole process to perturb a sample input:
@@ Appendix E More Details on Perturbation Method
Appendix F More Details about Training of Diffusion Models
In this section, we provide a detailed explanation of the fine-tuning methods used. We fine-tuned SD 2.1 with the best hyperparameters reported in the original papers for $\text{SFT}_{\text{Chosen}}$, Diffusion-DPO, and MaPO, using 8 A100×80 GB GPUs for all models. To fine-tune SD 2.1 with Diffusion and MaPO methods, we used a dataset $D=\{c,x^{w},x^{l}\}$ where $c,x^{w},x^{l}$ represent the prompt, preferred image, and less preferred image. To optimize a SD2.1 with $\text{SFT}_{\text{Chosen}}$ we utilized a dataset $D=\{c,x^{w}\}$ where $c,x^{w}$ represent the prompt, preferred image and image. In this paper, dataset $D$ represents the sampled and cleaned version of the Pick-a-Pic v2 dataset. Additionally, we clarify the DCPO models DCPO-c, DCPO-p, and DCPO-h. In this paper, DCPO-c and DCPO-p refer to SD 2.1 models fine-tuned with the DCPO method, using LLaVA and Emu2 for captioning and perturbation methods at three distinct levels, respectively. The main results for DCPO-p in the text are based on weak perturbation applied to the original prompt. In Table 3, we also report DCPO-p’s performance across other perturbation levels.
For DCPO-h, we applied perturbations to both the preferred and less preferred captions generated by LLaVA. The reported results for DCPO-h reflect a medium level of perturbation applied to the less preferred caption. In Table 3, we present the performance of DCPO-h across various perturbation levels, including perturbations to the preferred captions. Additionally, in Table 8, we show the results for DCPO-h using captions generated by Emu2.
Table 8: Results of the perturbation method applied to Emu2 captions across different levels.
DCPO-h ($z^{w},z^{w}_{p}$) weak 20.10 21.23 49.7 26.87 0.5003
DCPO-h ($z^{w},z^{l}_{p}$) weak 20.32 23.4 53.8 27.06 0.5070
DCPO-h ($z^{w},z^{w}_{p}$) medium 20.31 23.08 53.2 27.01 0.4895
DCPO-h ($z^{w},z^{l}_{p}$) medium 20.33 23.22 53.8 27.09 0.5009
DCPO-h ($z^{w},z^{w}_{p}$) strong 20.31 22.95 53.1 27.11 0.4878
DCPO-h ($z^{w},z^{l}_{p}$) strong 20.35 23.24 53.63 27.08 0.5050
The key findings indicate that perturbation on short captions not only fails to improve performance but also produces worse outcomes compared to DCPO-c (Emu2).
@@ Appendix F More Details about Training of Diffusion Models
Additionally, we conducted more experiments on in-distribution and out-of-distribution data. For this, we generated out-of-distribution data using LLaVA and Emu2 in the captioning setup. As shown in Figure 11, in-distribution data generally outperformed out-of-distribution data. However, the most significant improvement was observed with the hybrid method, as reported in Figure 6.
@@
Table 9 presents the performance details for different values of $\beta$, conducted using the medium level of DCPO-h. The results indicate that while lower values of $\beta$ significantly improve GenEval and HPSv2.1 on average, the optimal value for $\beta$ is 5000. We suggest that this hyperparameter may vary based on the dataset and task.
@@
DCPO-h 2500 20.53 25.81 58.0 27.02 0.5036
@@
Appendix G GPT-4o as an Evaluator
To obtain binary preferences from the API evaluator, we followed the approach outlined in the MaPO paper (Hong et al., 2024). Similar to Diffusion-DPO, we used three distinct questions to evaluate the images generated by the DCPO-h and Diffusion-DPO models, both utilizing SD 2.1 as the backbone. These questions were presented to the GPT-4o model to identify the preferred image. Below, we provide details of the prompts used.

### v2

@@ Acknowledgments
[…]
Appendix A Formalized Proofs of DCPO
A.1 Optimizing the DCPO Loss is Optimizing the DPO Loss
@@
[…]
@@
A.2 Analyses of DCPO’s Effectiveness
@@
$\Delta R_{\theta}\gg\Delta R_{\text{ref}}\implies\Delta_{\text{preferred}}\text{ increases.}$
@@
$\text{Relative Rate}=\frac{\Delta R_{\theta}}{\Delta R_{\text{ref}}}\gg 1,$
@@
Appendix B Supplementary Experiments and Analyses on DCPO
B.1 Semantic Overlap Distribution Measured in VQAScore
Figure 9: The conflict distribution that exists in the Pick-a-Pic V2 dataset in terms of VQAScore.
We also investigated the conflict distribution challenge of preference optimization datasets described in Section 3.1 using more recent prompt-image alignment measures, such as VQAScore (Karthik et al., 2023; Huang et al., 2023). Likewise, we calculated the VQAScores between each prompt $c$ and its preferred image, as well as between each prompt and its less-preferred image, from the Pick-a-Pic V2 dataset. Our results in Figure 9 indicate that, similar to the counterpart in terms of CLIPScore in Figure 3, there exists a consistently significant conflict between the semantic distributions of the preferred and less-preferred images with respect to the prompt $c$.
B.2 Comparison with Diffusion-KTO
@@
B.3 Comparison with Online and Interactive Methods
Beyond offline preference optimization methods such as Diffusion-DPO, MaPO, and Diffusion-KTO, we also evaluate iterative algorithms (e.g., SePPO) and online algorithms (e.g., SPO). Under a fixed training budget of 2,000 steps and a dataset of 20,000 samples, the results in Table 6 show that DCPO outperforms both SPO and SePPO, indicating greater cost efficiency and stronger alignment with fewer optimization steps and limited data. We leave a comprehensive comparison of offline and online methods under larger training budgets and datasets for future work.
Table 6: Comparison of DCPO-h against online (SPO) and iterative (SePPO) methods under a fixed training budget.
SPO 0.4449 19.30 24.40 46.45 25.91
SePPO (lr $1\times 10^{-8}$) 0.50567 20.34 25.09 55.21 26.83
SePPO (lr $5\times 10^{-9}$) 0.50232 20.32 25.08 55.24 26.82
DCPO-h 0.5100 20.57 25.62 58.20 27.13
B.4 Benchmarking Performance on Rapidata
To further demonstrate DCPO’s versatility, we fine-tune Stable Diffusion 2.1 using Diffusion-DPO and DCPO on another high-quality preference dataset Rapidata (Christodoulou & Kuhlmann-Jørgensen, 2024). Table 7 shows that our DCPO variants consistently outperform the Diffusion-DPO baseline on multiple benchmarking metrics, including GenEval, Pickscore, HPSv2.1, ImageReward, and CLIPscore.
Table 7: DCPO performances using SD2.1 on Rapidata (Christodoulou & Kuhlmann-Jørgensen, 2024).
Method (SD2.1) Geneval Pickscore HPSv2.1 ImageReward CLIPscore
Diffusion-DPO 0.4813 20.34 25.10 55.4 26.84
DCPO-c 0.4867 20.44 25.43 55.7 26.86
DCPO-h 0.4978 20.42 25.10 55.6 26.91
B.5 Performance of Using SDXL as the Backbone of DCPO
We also perform additional experiments to evaluate the performance of DCPO using Stable Diffusion XL (SDXL)  (Podell et al., 2023), a larger alternative backbone model for T2I instead of our default SD 2.1. Due to the larger parameter size of SDXL, we conduct LoRA fine-tuning with minimal hyperparameter search. The results in Table 8 show that DCPO outperforms Diffusion-DPO on key metrics such as Pickscore, Geneval, HPSv2, and CLIPscore, while achieving comparable performance on ImageReward. We hope that our results would encourage other researchers to further explore DCPO’s effectiveness on different datasets in future, gaining more valuable insights into its broader applicability and robustness.
Table 8: DCPO performances of using Stable Diffusion XL (SDXL) (Podell et al., 2023) as the backbone model.
Method (SDXL) Geneval (Overall) Pickscore HPSv2.1 ImageReward CLIPscore
Diffusion-DPO 0.5645 21.77 28.64 71.2 28.61
DCPO-c 0.5758 21.87 28.65 71.2 28.63
DCPO-h-weak 0.5704 21.87 28.64 71.2 28.62
DCPO-h-medium 0.5700 21.86 28.64 71.2 28.63
DCPO-h-strong 0.5696 21.86 28.64 71.2 28.62
Appendix C Pick-Double Caption Dataset
@@
Figure 10: The distribution of token lengths in the original prompts.
We observed that some prompts contained only one or two words, while others were excessively long. To ensure a fair comparison, we removed prompts that were too short or too long, leaving us with approximately 17,000 instances. We then generated captions using two state-of-the-art models, LLaVA-1.6-34B, and Emu2-32B. The construction of the Pick-Double Caption dataset is illustrated in Figure 11, which provides several examples.
We acknowledge that while captioning and perturbation introduce additional computational costs, these are one-time expenses incurred only during pre-processing and do not affect training or evaluation time. We introduce three variants of DCPO—c, p, and h—each with different processing requirements. For instance, DCPO-c and DCPO-h involve captioning, whereas DCPO-p is more computationally efficient as it only perturbs a less preferred caption. Furthermore, captioning 20,000 images using LLaVA requires less than 12 hours on a single A100:80G GPU, and this process can be significantly accelerated using multiple GPUs.
@@
Table LABEL:tab:tokens-detail presents a statistical analysis of the Pick-Double Caption dataset. With the non-conditional prompt method, we found that the average token length of captions generated by LLaVA is similar to that of the original prompts. However, captions generated by LLaVA using conditional prompts are twice as long as the original prompts. Additionally, Emu2 generated captions that, on average, are half the length of the original prompts for both methods.
@@
Appendix D Specifications of Creating Perturbed Captions
We provide the setups for the LLM-based perturbation process involved in the DCPO-p and DCPO-h pipelines. Similarly to the method of constructing paraphrasing adversarial attacks as synonym-swapping perturbation by Krishna et al. (2024), we use DIPPER, a text generation model built by fine-tuning T5-XXL (Chung et al., 2022), to create semantically perturbed captions or prompts, as shown in Table LABEL:tab:perturbation-examples. Our three levels of perturbation are achieved by only altering the setting of lexicon diversity (0 to 100) in DIPPER - we use 40 for Weak, 60 for Medium, and 80 for Strong. We also use "Text perturbation for variable text-to-image prompt." to prompt the perturbation. We hereby provide a code snippet to showcase the whole process to perturb a sample input as in Figure 12.
@@ Appendix D Specifications of Creating Perturbed Captions
Figure 12: Sample source code to create multiple perturbations (modified prompts) using DIPPER.
Appendix E Specifications of Perturbation Quality Control Using MLLM-as-a-Judge
We provide the evaluation prompt that defines the rubrics used for perturbation quality assessment, ranked by MLLM judges, as follows.
Evaluation Prompt for Assessing Perturbation (Modified Prompt) Quality Using an MLLM Judge You are given an image, and a text which is an image caption generated by another AI model for the given image. Rate how well the caption describes the image on a discrete scale from 1 to 5. Here is the detailed scoring rubric for evaluating the quality of the caption: Excellent (5): the caption perfectly adheres to the image, excelling in relevance, accuracy, comprehensiveness, creativity, and granularity. It provides an insightful, detailed, and thorough answer, indicating a deep and nuanced understanding of the image. Good (4): the caption is well-aligned with the image, demonstrating a high degree of relevance, accuracy, and comprehensiveness. It shows creativity and a nuanced understanding of the topic, with detailed granularity that enhances the response quality. Average (3): the caption adequately addresses the image, showing a fair level of relevance, accuracy, and comprehensiveness. It reflects a basic level of creativity and granularity but may lack sophistication or depth in fully capturing the image. Fair (2): the caption addresses the user’s instruction partially, with evident shortcomings in relevance, accuracy, or comprehensiveness. It lacks depth in creativity and granularity, indicating a superficial understanding of the image. Poor (1): the caption significantly deviates from the image and fails to address the query effectively. It shows a lack of relevance, accuracy, and comprehensiveness. Creativity and granularity are absent or poorly executed.
Per input, we ensure that any perturbation candidate does not outrank the input (i.e., the original prompts or the captions generated by LLaVa). For example, if the input is ranked as Good (4), its corresponding perturbations would be selected only from those that are ranked as Average (3) or below.
Appendix F Specifications of Model Fine-tuning
In this section, we detail the fine-tuning methods and hyperparameter searches conducted under a unified SD 2.1 setup using eight A100 80 GB GPUs. We fine-tune SD 2.1 for 2,000 training steps with each method. The dataset $D$ is a sampled and cleaned subset of Pick-a-Pic v2; each record comprises a text prompt $c$ and two images, where $x^{w}$ denotes the annotator-preferred image for $c$ and $x^{l}$ the less-preferred image.
$\text{SFT}_{\text{Chosen}}$.
We fine-tune SD 2.1 on pairs $\{c,x^{w}\}$. We sweep batch size in $\{64,128\}$ and learning rate in $\{10^{-10},10^{-9},10^{-8},10^{-7}\}$. The strongest configuration uses batch size $64$ with learning rate $10^{-9}$. Larger learning rates substantially degrade quality, and increasing the batch size to $128$ at fixed learning rate $10^{-9}$ reduces overall performance.
Diffusion-DPO.
We train SD 2.1 on triples $\{c,x^{w},x^{l}\}$ with batch size fixed to $128$ and learning rate fixed to $10^{-8}$. We sweep the regularization coefficient $\beta$ over $\{500,1000,2000,2500,5000\}$. Performance peaks at a moderate $\beta$ around $2000$, with small but consistent declines away from this setting, which aligns with the expected trade-off governed by $\beta$.
MaPO.
We also use triples $\{c,x^{w},x^{l}\}$ with batch size fixed to $128$ to fine-tune SD 2.1. Guided by the design of the method, we focus the search on the small-$\beta$ regime with $\beta\in\{0.01,0.1\}$ and learning rate in $\{10^{-8},10^{-7}\}$. The most favorable setting is $\beta=0.01$ with learning rate $10^{-8}$. Increasing $\beta$ or the learning rate degrades performance, reflecting sensitivity to heavier margin weighting and aggressive steps.
DCPO.
We explore two families of configurations that differ only in how captions and perturbations are constructed while keeping the SD 2.1 backbone and compute budget unchanged. DCPO-c and DCPO-p fine-tune SD 2.1 with the DCPO objective using captions produced by LLaVA and Emu2, respectively, and evaluate three perturbation levels. Unless otherwise noted, the main results for DCPO-p use a weak perturbation applied to the original prompt, and Table 3 reports DCPO-p performance across the other perturbation levels. DCPO-h uses distinct captions for each image; for every training instance we form two caption–image pairs, $(c^{w},x^{w})$ and $(c^{l},x^{l})$, where $c^{w}$ describes the preferred image and $c^{l}$ describes the less-preferred image. Captions are generated with LLaVA and then perturbed according to the DCPO-h protocol. For DCPO and Diffusion-DPO we sweep $\beta$ over $\{500,1000,2000,2500,5000\}$ with batch size $128$ and learning rate $10^{-8}$. Within this range DCPO attains its best overall performance at $\beta=5000$, and the reported DCPO-h results correspond to a medium perturbation applied to $c^{l}$. Table 3 presents DCPO-h performance across additional perturbation levels, including experiments that perturb $c^{w}$, and Table LABEL:tab:emu2-perturbed-caption-results reports DCPO-h results when captions are generated by Emu2. The key findings indicate that perturbations applied to short captions do not improve performance and often produce worse outcomes compared to DCPO-c (Emu2).
@@ Appendix F Specifications of Model Fine-tuning
We also examine sensitivity to distribution shift by constructing in-distribution and out-of-distribution splits using LLaVA and Emu2 within the captioning setup. As shown in Figure 13, in-distribution data generally outperforms out-of-distribution data, while the most significant improvement is observed with the hybrid configuration reported in Figure 6.
Finally, Table 9 reports the $\beta$ sweep for DCPO-h under the medium perturbation setting alongside Diffusion-DPO. The results highlight that, across a range of $\beta$ values, DCPO-h can outperform Diffusion-DPO, underscoring the effectiveness of the DCPO framework. Based on these observations, we recommend validating $\beta$ for each dataset and task, since the preferred setting can shift with distributional differences and metric emphasis.
@@
Diffusion-DPO 1000 20.34 24.99 55.06 26.86 0.4935
DCPO-h 2000 20.61 25.83 58.36 27.19 0.5056
Diffusion-DPO 2000 20.36 25.10 56.43 26.98 0.4858
DCPO-h 2500 20.53 25.81 58.00 27.02 0.5036
Diffusion-DPO 2500 20.35 25.11 55.24 26.86 0.4920
@@
Diffusion-DPO 5000 20.34 25.16 55.49 26.87 0.4974
Appendix G GPT-4o as Evaluator
Positional Bias.
To obtain binary preferences from the API evaluator, we follow the approach outlined in the MaPO paper (Hong et al., 2024). To address positional bias in GPT-4o’s evaluations, we alternate the positions of the images across different criteria on 100 samples (explained in Section 4). The results in Table 10 show that DCPO consistently achieves better performance than Diffusion-DPO, even when positional bias is accounted for.
Table 10: Comparison of DCPO and Diffusion-DPO across different positions in prompt of GPT-4o based on various criteria.
Method General Preference Visual Appeal Prompt Alignment
DCPO-h 58% 64.5% 56.5%
Diffusion-DPO 42% 35.5% 43.5%
Human Study.
While recent work Hong et al. (2024) shows that GPT-4o can judge two images given the same prompt, we conducted a small human study to independently assess this alignment. We randomly sampled 100 image pairs produced by SD 2.1 fine-tuned with DCPO-h and Diffusion-DPO, and recruited three volunteer researchers to evaluate visual appeal. Each annotator answered the question “Which image is more visually appealing?” The comparison performance of DCPO-h and Diffusion-DPO on Visual Appeal, reported in Tables 10 and 11, indicates that GPT-4o judgments closely align with human preferences.
Table 11: Comparison of DCPO and Diffusion-DPO across different positions in the prompt of GPT-4o based on various criteria.
Method Volunteer 1 Volunteer 2 Volunteer 3 Average
Win Rate of DCPO-h 64% 66% 69% 66.33%
Win Rate of Diffusion-DPO 36% 34% 41% 33.67%
Judgment Prompts.
Similarly to Diffusion-DPO, we use three distinct questions to evaluate the images generated by DCPO-h and other baseline models, all utilizing SD 2.1 as the backbone. These questions are presented to GPT-4o to identify the preferred image. In the following, we provide details of the prompts used.